OpenStudy (anonymous):

During the winter, an ice cream shop has seventy-five percent of the employees it has during the summer. Which equation shows that during the winter, the ice cream shop has twelve employees? 12=0.75*s 75=0.12*s w=0.75*12 w=12%*75

OpenStudy (nincompoop):

do you know what a per cent is?

OpenStudy (anonymous):

yes yes i do:)

OpenStudy (nincompoop):

okay then answer the question

OpenStudy (nincompoop):

http://www.mathsisfun.com/percentage.html

OpenStudy (anonymous):

not to be rude but can i just get the answere bc i need to be done befor the 18

OpenStudy (nincompoop):

that would be cheating.

OpenStudy (nincompoop):

do you want to be labeled as the cheater?

OpenStudy (anonymous):

not really its just helping someone go to 8th grade :)

OpenStudy (nincompoop):

I will give you the answer, but you need to tell your teacher and your parents that you cheated

OpenStudy (nincompoop):

I want them to sign a paper saying that you've notified them of your cheating habits

OpenStudy (anonymous):

ok sure:)

OpenStudy (nincompoop):

come back when you have the document

OpenStudy (anonymous):

answere plese :)

OpenStudy (anonymous):

ill give u a medal :)

OpenStudy (here_to_help15):

i will :P

OpenStudy (anonymous):

ok!!

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Please don't give him the answer.. show them how to do it and let them work it out for themselves

OpenStudy (anonymous):

ok explaine:)

OpenStudy (ash2326):

@AntonioCochrane Could you tell me the relation between employees in winter to employees in summer?

OpenStudy (anonymous):

um.........:)

OpenStudy (anonymous):

the both worck

OpenStudy (here_to_help15):

C) w = ( 0.75 ) ( S ) = ( 0.75 )( 12 ) = 9 winter employes

OpenStudy (anonymous):

ok ok

OpenStudy (here_to_help15):

itz c :P

OpenStudy (anonymous):

Good try ash

OpenStudy (anonymous):

u should have shown me how to do it not give me answere. but thanks:)

OpenStudy (ash2326):

@AntonioCochrane Wait

OpenStudy (here_to_help15):

srry :(

OpenStudy (anonymous):

i dont want the answere ill do it on my own :)

OpenStudy (ash2326):

This is not the correct answer, it says that Employees in winter are 12 We don;t have to find the no. of employees but the equation which will help in solving

OpenStudy (anonymous):

Dude, you just asked for the answer, though.

OpenStudy (anonymous):

ik ill just do it on my own

OpenStudy (anonymous):

K

OpenStudy (anonymous):

ok:)

OpenStudy (nincompoop):

seventy-five percent of the employees it has during the summer W = 75% S equation shows that during the winter, the ice cream shop has twelve employees? W = 12 12 = 75% S percent, decimal and fraction are all related

OpenStudy (nincompoop):

let us try the concept of dividing pizza into several slices say a pan of pizza was divided into four slices and each slice will be represented as 1 of 4 is the same as 1/4 and 1/4 = 0.25 since 4*0.25 = 1.00 to convert decimal to per cent we need to first perform the indicated operation in fraction then multiply it by 100 (per cent, cent means 100) 1/4 * 100 = 0.25 * 100 = 25%

OpenStudy (anonymous):

ooo now u want to help me :p